# Bilayer Cuprate Antiferromagnets Enable Programmable Cavity Optomagnonics

### v1

@@ I Introduction
Engineered light-matter interactions enable transformative advances across classical photonics Wang et al. (2022); Parvini et al. (2017); Pontula et al. (2025); Badloe and Rho (2024); Winkel et al. (2024); Parvini and Khazaei Nezhad (2022); Parvini et al. (2015), neuromorphic computing architectures Tait et al. (2017); Li et al. (2025); Prucnal and Shastri (2017); Oberbauer et al. (2025), and quantum information technologies Chan et al. (2023); Sekine et al. (2024); Liu et al. (2025); Zhang (2023); Ying (2023); Shashank et al. (2025). Cavity magnonics—where magnons hybridize with microwave and optical photons—enables tunable quantum dynamics Hirayama et al. (2021); Song et al. (2025); Ghirri et al. (2023); Lee et al. (2023); Rani et al. (2025); Liu et al. (2016); Rameshti et al. (2022); Osada et al. (2018); Liang et al. (2023); Haigh et al. (2016); Šimić et al. (2020); Zhang et al. (2014); Hisatomi et al. (2016). Yttrium iron garnet (YIG) Kornack et al. (2007); Soykal and Flatté (2010a, b); Huebl et al. (2013); Tabuchi et al. (2014) established the benchmark with cooperativity exceeding $10^{5}$ Goryachev et al. (2014); Bourhill et al. (2016); Xu et al. (2024), multi-GHz splittings, multimode hybridization involving photons, magnons, and phonons Zuo et al. (2024); Shen et al. (2025); Tabuchi et al. (2016), and Gilbert damping $8.6\times 10^{-5}$ Guo et al. (2023), achieving single-magnon detection and quantum-limited sensing. YIG still faces engineering challenges, including substrate-driven trade-offs between mode volume and coherence, increased millikelvin damping from paramagnetic substrates, and a narrower accessible frequency range than antiferromagnets Li et al. (2020); Soykal and Flatté (2010b), motivating exploration of alternative magnetic platforms.
@@
Antiferromagnetic (AFM) insulators have transformed cavity magnonics through fundamental advantages over ferrimagnets: negligible stray fields, ultrafast GHz-THz dynamics, and robust frequency tunability Lachance-Quirion et al. (2019); Parvini et al. (2025, 2020); Frostad et al. (2025); Kaarbø et al. (2025); Falch et al. (2025a); Shiranzaei et al. (2023). Recent terahertz-cavity experiments with NiO have realized cavity magnon–polaritons in the THz band Kritzell et al. (2024); Moriyama et al. (2019), and Hematite ($\alpha$-Fe2O3) has shown room-temperature strong magnon–photon coupling, underscoring that AFMs can operate at high frequencies with low loss Białek et al. (2021); Boventer et al. (2023); Yang et al. (2022); Ghirri et al. (2023); Białek et al. (2023). Cuprate AFMs—exemplified by YBa2Cu3O6+x (YBCO) and Bi2Sr2CaCu2O8+x (BSCCO)—are paradigmatic correlated materials in which intertwined spin, charge, and lattice degrees of freedom support high-$T_{c}$ superconductivity and robust antiferromagnetism Tranquada et al. (1988); Singh et al. (1992); Ajay et al. (1995); Pratap et al. (1999); Jin et al. (2025); Bounoua et al. (2022); Reznik et al. (2006); Le Tacon et al. (2011); Tranquada (2007); Bao et al. (2025). Their spin dynamics are set by a remarkably large in-plane superexchange, $J_{\parallel}\!\sim\!130~\mathrm{meV}$, and quasi-two-dimensional correlations that persist well above $T_{N}$ Haug et al. (2010); Shamoto et al. (1993). These extreme energy scales, together with the bilayer structure of YBCO, make cuprates natural candidates for cavity magnonics and optomagnonics. Theory predicts that cuprates embedded in THz cavities can realize both linear and quadratic photon–magnon interactions, including a bilayer-specific bimagnon channel Curtis et al. (2022); Ghirri et al. (2015). These advances position cuprate AFMs as a promising cavity-QED platform.
@@ II Bilayer Cuprate AFM Hamiltonian
The parent compound YBa2Cu3O6+x (YBCO) exhibits long-range antiferromagnetic order arising from its bilayer CuO2 planes. Each unit cell hosts a CuO2 bilayer separated by yttrium (Fig. 1a), with in-plane lattice constant $a\simeq 3.86$ Å, $c$-axis parameter $c\simeq 11.7$ Å, and fractional bilayer offset $z=0.28$, giving a CuO2–CuO2 spacing $zc\simeq 3.3$ Å. Across the oxygenation range that retains long-range AFM order, the Néel temperature decreases with $x$. Near $x=0$ $T_{N}\approx 420$ K, at $x\approx 0.1$ $T_{N}\approx 350\pm 50$ K, at $x\approx 0.2$ $T_{N}\approx 250\pm 50$ K, and long-range AFM typically vanishes by $x\gtrsim 0.35$$0.40$. The dominant coupling is the AFM superexchange within a plane, $J_{\parallel}\sim(80\text{–}120)$ meV from neutron and Raman scattering Tranquada et al. (1989); Abragam and Bleaney (2012); Chen et al. (2011); Vettier et al. (1989); Chubukov and Frenkel (1995). The bilayer stacking introduces a direct intra-bilayer exchange $J_{\perp 1}\approx 2.6$ meV between Cu atoms in adjacent planes, while a much weaker inter-bilayer superexchange $J_{\perp 2}\approx 0.026$ meV couples successive bilayers through the Cu–O chain layer. This hierarchy $J_{\parallel}\gg J_{\perp 1}\gg J_{\perp 2}$ enforces quasi-two-dimensional correlations with limited three-dimensional coherence. Spin–orbit coupling induces a weak easy-plane anisotropy that suppresses out-of-plane fluctuations, confines the ordered moments to the $ab$ plane, and, for definiteness, sets the Néel vector along $\hat{x}$. We apply a static in-plane field along $\hat{\mathbf{y}}$, which induces a transverse canting within the easy plane. The spin-$\tfrac{1}{2}$ Hamiltonian is
@@
where $\mathbf{S}_{n}^{\kappa}$ denotes the spin operator at site $n$ in layer $\kappa\in\{1,2\}$ and $g_{\mathrm{Cu}}\simeq 2.15$ for Cu2+. The transverse field cants the spins by an angle $\chi$, giving the rotated-frame vector $\tilde{\mathbf{S}}$ Vishina et al. (2012); Boström et al. (2023):
@@
The classical ground-state energy per unit cell is $E_{\mathrm{cl}}/N=2g_{\mathrm{Cu}}\mu_{\mathrm{B}}B\sin\chi-(2J_{\parallel}+\overline{J}_{\perp})(1-2\sin^{2}\chi)$, and minimization with respect to $\chi$ gives the equilibrium canting angle Maksimov and Chernyshev (2016); Boström et al. (2021, 2023):
@@
here $B_{\mathrm{Z}}\equiv g_{\mathrm{Cu}}\mu_{\mathrm{B}}B$, at $B=0$ energies reduce to 0, $2J_{\parallel}\sqrt{2\alpha_{D}\mathcal{J}}$, $2J_{\parallel}\sqrt{(2-\alpha_{D})\,\overline{J}_{\perp}/J_{\parallel}}$, and $2J_{\parallel}\sqrt{2\!\left(\alpha_{D}+\overline{J}_{\perp}/J_{\parallel}\right)}$, respectively. The field dependence of the $\Gamma$-point magnon modes is shown in Fig. 1d. Most importantly, because $\alpha_{D}\ll\mathcal{J}$, the $\alpha$ mode is weakly influenced by anisotropy and exhibits nearly perfect linear Zeeman scaling, $E_{\alpha}\simeq B$, while remaining gapless at zero field. This hallmark behavior identifies it as the Goldstone mode associated with spontaneous U(1) symmetry breaking. Mode $\eta$ shows negligible anisotropy dependence and acquires only a weak quadratic correction under applied field. By contrast, the anisotropy-dominated $\beta$ and $\zeta$ modes scale as $\propto\sqrt{\alpha_{D}}$ and remain essentially field-independent. For reference, the Supplementary Information compares YIG and cuprate $\alpha$-mode frequencies. Despite higher damping in cuprates, the $\alpha$ mode shows a steeper Zeeman slope ($\sim$30 vs. 28 GHz/T) and coexists with THz optical branches, enabling broadband (GHz–THz) hybrid cavity architectures beyond YIG-only platforms.
In this work, we exploit the gapless $\alpha$ mode as a quantum gateway, coupling it to microwave cavities to open a new frontier in magnonic quantum electrodynamics. Its entanglement with the higher-frequency $\beta$ mode, rooted in their shared Bogoliubov origin Yuan et al. (2020); Mousolou et al. (2020), enables coherent transfer of quantum information between microwave and THz scales, effectively turning the bilayer cuprate into a solid-state frequency converter. For the bulk anisotropy $\alpha_{D}=1.0\times 10^{-4}$ Tranquada et al. (1989), the $\beta$ branch lies at $686~\mathrm{GHz}$, with the $\alpha$$\beta$ crossing at $B=22.8~\mathrm{T}$. At this degeneracy, magneto-optical probes such as MOKE can address the $\beta$ mode while microwave techniques control the $\alpha$ mode, realizing a microwave–THz quantum transducer spanning nearly three orders of magnitude in frequency. This functionality is tunable: reducing anisotropy via epitaxial strain, oxygen stoichiometry, or targeted substitution Vishina et al. (2012); Ihara et al. (1997); Burlet et al. (1998); Kochelaev (2016) lowers $f_{\beta}$ and shifts the crossing into experimentally accessible fields. For $\alpha_{D}=1.0\times 10^{-5}$, the modes resonate at $217~\mathrm{GHz}$ with $B=7.22~\mathrm{T}$, already compatible with emerging THz quantum technologies. Driving the anisotropy further down brings both branches into the microwave domain, enabling complete quantum protocols with standard cavity instrumentation. This tunability positions bilayer cuprates as a uniquely versatile platform, functioning either as a microwave–THz transducer or as an all-microwave architecture for scalable quantum information processing.
@@ III Optomagnonic Coupling
The quantized magnetic field in the cavity follows standard field quantization Yuan and Wang (2017); Johansen and Brataas (2018); Xiao et al. (2019):
@@
where $\hat{p}_{\lambda\mathbf{k}}^{\dagger}$ ($\hat{p}_{\lambda\mathbf{k}}$) creates (annihilates) a photon with wave vector $\mathbf{k}$, frequency $\omega_{\mathbf{k}}$, and helicity $\lambda=\pm 1$, and $V_{\mathrm{B}}$ is the effective cavity mode volume. The propagation direction is $\hat{\mathbf{k}}=\sin\theta\cos\phi\,\hat{\mathbf{x}}+\sin\theta\sin\phi\,\hat{\mathbf{y}}+\cos\theta\,\hat{\mathbf{z}}$ with circular polarization vector $\mathbf{e}_{\lambda}=(e_{x}\hat{\mathbf{x}}+e_{y}\hat{\mathbf{y}}+e_{z}\hat{\mathbf{z}})/\sqrt{2}$, where $e_{x}=-\sin\phi-\mathrm{i}\lambda\cos\theta\cos\phi$, $e_{y}=\cos\phi-\mathrm{i}\lambda\cos\theta\sin\phi$, and $e_{z}=\mathrm{i}\lambda\sin\theta$. Restricting to a single cavity mode at resonance frequency $\omega_{c}$ with $(\lambda_{0},\mathbf{k}_{0})$, the cavity Hamiltonian reduces to $\hat{\mathcal{H}}_{c}=\hbar\omega_{c}\big(\hat{p}^{\dagger}\hat{p}+1/2\big)$, where $\hat{p}\equiv\hat{p}_{\lambda_{0}\mathbf{k}_{0}}$. The spins couple to the cavity photon field via the Zeeman interaction $\hat{\mathcal{H}}_{\mathrm{I}}=g_{\mathrm{Cu}}\mu_{\mathrm{B}}\sum_{n}\hat{\mathbf{B}}_{\mathrm{p}}(\mathbf{r}_{n})\cdot\hat{\mathbf{S}}_{n}$. After Holstein–Primakoff and Fourier transformations, this yields (see SI)
@@
where the upper (lower) sign corresponds to $\Lambda_{1,3}$ ($\Lambda_{2,4}$). The collective coupling strength is $g_{0}=g_{\mathrm{Cu}}\mu_{\mathrm{B}}\sqrt{\mu_{0}\hbar\omega_{c}\rho\eta_{B}}$ with spin density $\rho=N/V_{\mathrm{s}}$ (for YBCO, $\rho\approx 1.1\times 10^{28}\,\mathrm{m}^{-3}$ Tranquada et al. (1989)) and filling factor $\eta_{B}=V_{\mathrm{s}}/V_{\mathrm{B}}\leq 1$ quantifying the overlap between the magnetic sample and cavity field Huebl et al. (2013).
@@
revealing the cavity as a tunable mediator of inter-magnon exchange. Reducing the cavity–magnon detuning $(\omega_{m}-\omega_{c})$—by tuning $f_{c}$ at the $\alpha$$\beta$ degeneracy—enhances the cavity-mediated exchange $(\Delta f_{\alpha\beta})$ while $C_{\alpha\beta}\!\ll\!1$ keeps the dynamics dispersive, providing a single practical knob for otherwise negligible inter-magnon coupling, unlike platforms requiring multiple magnets or engineered degeneracies Li et al. (2022); Xu et al. (2019).
@@
showing that the cavity couples only to the bright mode $\hat{m}$ with collective coupling energy $\mathcal{G}$, while the dark mode $\hat{n}$ is entirely decoupled. The resulting spectrum consists of two bright polaritons at $f_{0}\pm\mathcal{G}/h$ and a dark magnon at $f_{0}$. Thus, at triple resonance the photon–magnon and magnon–magnon gaps reduce to a single characteristic frequency scale $\mathcal{G}/h$: the bright–bright separation is $2\,\mathcal{G}/h$, and the nearest-neighbor gaps are each $\mathcal{G}/h$, explaining the convergence of the two curves in Fig. 2(c,d). This three-mode reorganization into bright and dark superpositions is a distinctive feature of bilayer cuprates, beyond the conventional two-mode Rabi doublet of ferrimagnets. The convergence of photon–magnon and magnon–magnon couplings into a single collective scale provides a powerful tuning knob, while the dark magnon—decoupled from cavity dissipation—offers long-lived storage, loss-free coupling, and noise resilience. The bright mode supplies a strong interface for signal conversion, establishing bilayer cuprates as a natural platform for memory and transduction Gardin et al. (2024); Zhao et al. (2023); Falch et al. (2025b).
The parameter dependence in Fig. 2(c,d) highlights two clear tuning knobs. Increasing the magnetic filling factor $\rho\eta_{\mathrm{B}}$ boosts both $g_{\alpha p}$ and the cavity-mediated $g_{\alpha\beta}$, enhancing the overall coupling strength. By contrast, increasing the anisotropy shifts the triple resonance to higher frequencies and fields. Together, these controls provide a direct way to balance photon–magnon and magnon–magnon interactions in the cavity spectrum. Because $\Gamma$-point AFMR linewidths in cuprates are not reliably established, we adopt a conservative effective linewidth of $\gamma_{\alpha}=0.2$ GHz and $\gamma_{\beta}=0.5$ GHz Wang and Xiao (2024).

### v2

@@ I Introduction
Engineered light-matter interactions enable transformative advances across classical photonics [1, 2, 3, 4, 5, 6, 7], neuromorphic computing architectures [8, 9, 10, 11, 12], and quantum information technologies [13, 14, 15, 16, 17, 18]. Cavity magnonics—where magnons hybridize with microwave and optical photons—enables tunable quantum dynamics [19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Yttrium iron garnet (YIG) [32, 33, 34, 35, 36] established the benchmark with cooperativity exceeding $10^{5}$ [37, 38, 39], multi-GHz splittings, multimode hybridization involving photons, magnons, and phonons [40, 41, 42], and Gilbert damping $8.6\times 10^{-5}$ [43], achieving single-magnon detection and quantum-limited sensing. YIG still faces engineering challenges, including substrate-driven trade-offs between mode volume and coherence, increased millikelvin damping from paramagnetic substrates, and a narrower accessible frequency range than antiferromagnets [44, 34], motivating exploration of alternative magnetic platforms.
@@
Antiferromagnetic (AFM) insulators have transformed cavity magnonics through fundamental advantages over ferrimagnets: negligible stray fields, ultrafast GHz-THz dynamics, and robust frequency tunability [45, 46, 47, 48, 49, 50, 51]. Recent terahertz-cavity experiments with NiO have realized cavity magnon–polaritons in the THz band [52, 53], and Hematite ($\alpha$-Fe2O3) has shown room-temperature strong magnon–photon coupling, underscoring that AFMs can operate at high frequencies with low loss [54, 55, 56, 21, 57]. Cuprate AFMs—exemplified by YBa2Cu3O6+x (YBCO) and Bi2Sr2CaCu2O8+x (BSCCO)—are paradigmatic correlated materials in which intertwined spin, charge, and lattice degrees of freedom support high-$T_{c}$ superconductivity and robust antiferromagnetism [58, 59, 60, 61, 62, 63, 64, 65, 66, 67]. Their spin dynamics are set by a remarkably large in-plane superexchange, $J_{\parallel}\!\sim\!130~\mathrm{meV}$, and quasi-two-dimensional correlations that persist well above $T_{N}$ [68, 69]. These extreme energy scales, together with the bilayer structure of YBCO, make cuprates natural candidates for cavity magnonics and optomagnonics. Theory predicts that cuprates embedded in THz cavities can realize both linear and quadratic photon–magnon interactions, including a bilayer-specific bimagnon channel [70, 71]. These advances position cuprate AFMs as a promising cavity-QED platform.
@@ II Bilayer Cuprate AFM Hamiltonian
The parent compound YBa2Cu3O6+x (YBCO) exhibits long-range antiferromagnetic order arising from its bilayer CuO2 planes. Each unit cell hosts a CuO2 bilayer separated by yttrium (Fig. 1a), with in-plane lattice constant $a\simeq 3.86$ Å, $c$-axis parameter $c\simeq 11.7$ Å, and fractional bilayer offset $z=0.28$, giving a CuO2–CuO2 spacing $zc\simeq 3.3$ Å. Across the oxygenation range that retains long-range AFM order, the Néel temperature decreases with $x$. Near $x=0$ $T_{N}\approx 420$ K, at $x\approx 0.1$ $T_{N}\approx 350\pm 50$ K, at $x\approx 0.2$ $T_{N}\approx 250\pm 50$ K, and long-range AFM typically vanishes by $x\gtrsim 0.35$$0.40$. The dominant coupling is the AFM superexchange within a plane, $J_{\parallel}\sim(80\text{–}120)$ meV from neutron and Raman scattering [72, 73, 74, 75, 76]. The bilayer stacking introduces a direct intra-bilayer exchange $J_{\perp 1}\approx 2.6$ meV between Cu atoms in adjacent planes, while a much weaker inter-bilayer superexchange $J_{\perp 2}\approx 0.026$ meV couples successive bilayers through the Cu–O chain layer. This hierarchy $J_{\parallel}\gg J_{\perp 1}\gg J_{\perp 2}$ enforces quasi-two-dimensional correlations with limited three-dimensional coherence. Spin–orbit coupling induces a weak easy-plane anisotropy that suppresses out-of-plane fluctuations, confines the ordered moments to the $ab$ plane, and, for definiteness, sets the Néel vector along $\hat{x}$. We apply a static in-plane field along $\hat{\mathbf{y}}$, which induces a transverse canting within the easy plane. The spin-$\tfrac{1}{2}$ Hamiltonian is
@@
where $\mathbf{S}_{n}^{\kappa}$ denotes the spin operator at site $n$ in layer $\kappa\in\{1,2\}$ and $g_{\mathrm{Cu}}\simeq 2.15$ for Cu2+. The transverse field cants the spins by an angle $\chi$, giving the rotated-frame vector $\tilde{\mathbf{S}}$ [77, 78]:
@@
The classical ground-state energy per unit cell is $E_{\mathrm{cl}}/N=2g_{\mathrm{Cu}}\mu_{\mathrm{B}}B\sin\chi-(2J_{\parallel}+\overline{J}_{\perp})(1-2\sin^{2}\chi)$, and minimization with respect to $\chi$ gives the equilibrium canting angle [79, 80, 78]:
@@
here $B_{\mathrm{Z}}\equiv g_{\mathrm{Cu}}\mu_{\mathrm{B}}B$, at $B=0$ energies reduce to 0, $2J_{\parallel}\sqrt{2\alpha_{D}\mathcal{J}}$, $2J_{\parallel}\sqrt{(2-\alpha_{D})\,\overline{J}_{\perp}/J_{\parallel}}$, and $2J_{\parallel}\sqrt{2\!\left(\alpha_{D}+\overline{J}_{\perp}/J_{\parallel}\right)}$, respectively. The field dependence of the $\Gamma$-point magnon modes is shown in Fig. 1d. Most importantly, because $\alpha_{D}\ll\mathcal{J}$, the $\alpha$ mode is weakly influenced by anisotropy and exhibits nearly perfect linear Zeeman scaling, $E_{\alpha}\simeq B$, while remaining gapless at zero field. This hallmark behavior identifies it as the Goldstone mode associated with spontaneous U(1) symmetry breaking. Mode $\eta$ shows negligible anisotropy dependence and acquires only a weak quadratic correction under applied field. By contrast, the anisotropy-dominated $\beta$ and $\zeta$ modes scale as $\propto\sqrt{\alpha_{D}}$ and remain essentially field-independent. For reference, the Supplementary Information compares YIG and cuprate $\alpha$-mode frequencies. Despite higher damping in cuprates, the $\alpha$ mode shows a steeper Zeeman slope ($\sim$30 vs. 28 GHz/T) and coexists with THz optical branches, enabling broadband (GHz–THz) hybrid cavity architectures beyond YIG-only platforms [81].
In this work, we exploit the gapless $\alpha$ mode as a quantum gateway, coupling it to microwave cavities to open a new frontier in magnonic quantum electrodynamics. Its entanglement with the higher-frequency $\beta$ mode, rooted in their shared Bogoliubov origin [82, 83], enables coherent transfer of quantum information between microwave and THz scales, effectively turning the bilayer cuprate into a solid-state frequency converter. For the bulk anisotropy $\alpha_{D}=1.0\times 10^{-4}$ [72], the $\beta$ branch lies at $686~\mathrm{GHz}$, with the $\alpha$$\beta$ crossing at $B=22.8~\mathrm{T}$. At this degeneracy, magneto-optical probes such as MOKE can address the $\beta$ mode while microwave techniques control the $\alpha$ mode, realizing a microwave–THz quantum transducer spanning nearly three orders of magnitude in frequency. This functionality is tunable: reducing anisotropy via epitaxial strain, oxygen stoichiometry, or targeted substitution [77, 84, 85, 86] lowers $f_{\beta}$ and shifts the crossing into experimentally accessible fields. For $\alpha_{D}=1.0\times 10^{-5}$, the modes resonate at $217~\mathrm{GHz}$ with $B=7.22~\mathrm{T}$, already compatible with emerging THz quantum technologies. Driving the anisotropy further down brings both branches into the microwave domain, enabling complete quantum protocols with standard cavity instrumentation. This tunability positions bilayer cuprates as a uniquely versatile platform, functioning either as a microwave–THz transducer or as an all-microwave architecture for scalable quantum information processing.
@@ III Optomagnonic Coupling
The quantized magnetic field in the cavity follows standard field quantization [87, 88, 89]:
@@
where $\hat{p}_{\lambda\mathbf{k}}^{\dagger}$ ($\hat{p}_{\lambda\mathbf{k}}$) creates (annihilates) a photon in the mode characterized by wave vector $\mathbf{k}$, frequency $\omega_{\mathbf{k}}$, and helicity $\lambda=\pm 1$, and $V_{\mathrm{B}}$ is the effective cavity mode volume. The propagation direction is $\hat{\mathbf{k}}=\sin\theta\cos\phi\,\hat{\mathbf{x}}+\sin\theta\sin\phi\,\hat{\mathbf{y}}+\cos\theta\,\hat{\mathbf{z}}$ with circular polarization vector $\mathbf{e}_{\lambda}=(e_{x}\hat{\mathbf{x}}+e_{y}\hat{\mathbf{y}}+e_{z}\hat{\mathbf{z}})/\sqrt{2}$, where $e_{x}=-\sin\phi-\mathrm{i}\lambda\cos\theta\cos\phi$, $e_{y}=\cos\phi-\mathrm{i}\lambda\cos\theta\sin\phi$, and $e_{z}=\mathrm{i}\lambda\sin\theta$. Restricting to a single cavity mode at resonance frequency $\omega_{c}$ with $(\lambda_{0},\mathbf{k}_{0})$, the cavity Hamiltonian reduces to $\hat{\mathcal{H}}_{c}=\hbar\omega_{c}\big(\hat{p}^{\dagger}\hat{p}+1/2\big)$, where $\hat{p}\equiv\hat{p}_{\lambda_{0}\mathbf{k}_{0}}$. The spins couple to the cavity photon field via the Zeeman interaction $\hat{\mathcal{H}}_{\mathrm{I}}=g_{\mathrm{Cu}}\mu_{\mathrm{B}}\sum_{n}\hat{\mathbf{B}}_{\mathrm{p}}(\mathbf{r}_{n})\cdot\hat{\mathbf{S}}_{n}$. After Holstein–Primakoff and Fourier transformations, this yields (see SI)
@@
where the upper (lower) sign corresponds to $\Lambda_{1,3}$ ($\Lambda_{2,4}$). The collective coupling strength is $g_{0}=g_{\mathrm{Cu}}\mu_{\mathrm{B}}\sqrt{\mu_{0}\hbar\omega_{c}\rho\eta_{B}}$ with spin density $\rho=N/V_{\mathrm{s}}$ (for YBCO, $\rho\approx 1.1\times 10^{28}\,\mathrm{m}^{-3}$ [72]) and filling factor $\eta_{B}=V_{\mathrm{s}}/V_{\mathrm{B}}\leq 1$ quantifying the overlap between the magnetic sample and cavity field [35].
@@
revealing the cavity as a tunable mediator of inter-magnon exchange. Reducing the cavity–magnon detuning $(\omega_{m}-\omega_{c})$—by tuning $f_{c}$ at the $\alpha$$\beta$ degeneracy—enhances the cavity-mediated exchange $(\Delta f_{\alpha\beta})$ while $C_{\alpha\beta}\!\ll\!1$ keeps the dynamics dispersive, providing a single practical knob for otherwise negligible inter-magnon coupling, unlike platforms requiring multiple magnets or engineered degeneracies [92, 90].
@@
showing that the cavity couples only to the bright mode $\hat{m}$ with collective coupling energy $\mathcal{G}$, while the dark mode $\hat{n}$ is entirely decoupled. The resulting spectrum consists of two bright polaritons at $f_{0}\pm\mathcal{G}/h$ and a dark magnon at $f_{0}$. Thus, at triple resonance the photon–magnon and magnon–magnon gaps reduce to a single characteristic frequency scale $\mathcal{G}/h$: the bright–bright separation is $2\,\mathcal{G}/h$, and the nearest-neighbor gaps are each $\mathcal{G}/h$, explaining the convergence of the two curves in Fig. 2(c,d). This three-mode reorganization into bright and dark superpositions is a distinctive feature of bilayer cuprates, beyond the conventional two-mode Rabi doublet of ferrimagnets. The convergence of photon–magnon and magnon–magnon couplings into a single collective scale provides a powerful tuning knob, while the dark magnon—decoupled from cavity dissipation—offers long-lived storage, loss-free coupling, and noise resilience. The bright mode supplies a strong interface for signal conversion, establishing bilayer cuprates as a natural platform for memory and transduction [93, 94, 95].
The parameter dependence in Fig. 2(c,d) highlights two clear tuning knobs. Increasing the magnetic filling factor $\rho\eta_{\mathrm{B}}$ boosts both $g_{\alpha p}$ and the cavity-mediated $g_{\alpha\beta}$, enhancing the overall coupling strength. By contrast, increasing the anisotropy shifts the triple resonance to higher frequencies and fields. Together, these controls provide a direct way to balance photon–magnon and magnon–magnon interactions in the cavity spectrum. Because $\Gamma$-point AFMR linewidths in cuprates are not reliably established, we adopt a conservative effective linewidth of $\gamma_{\alpha}=0.2$ GHz and $\gamma_{\beta}=0.5$ GHz [96].